OpenStudy (kittiwitti1):

Graph the unit circle using parametric equations with your calculator set to degree mode. Use a scale of 5. Trace the circle to find all values of t between 0° and 360° satisfying the following statement. (Enter your answers as a comma-separated list.) A) sin t = cos t B) −sin t = cos t If angle θ is in standard position and the terminal side of θ intersects the unit circle at the point \[\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}}\]find csc θ, sec θ, and cot θ.

OpenStudy (anonymous):

The equation you did, didn't how correctly

OpenStudy (kittiwitti1):

What equation ... I didn't do any equation

OpenStudy (anonymous):

\[\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}}\] whats that

OpenStudy (anonymous):

nvm

OpenStudy (kittiwitti1):

That's what the problem says. Is that wrong?

OpenStudy (kittiwitti1):

lol

OpenStudy (anonymous):

it shows differently on my screen

Directrix (directrix):

Mine too. See attachment.

OpenStudy (anonymous):

SEE

OpenStudy (unicornpoopcookies):

That question seems to already have been asked on here

OpenStudy (unicornpoopcookies):

check this maybe hang on

OpenStudy (kittiwitti1):

Oh. Well here you go http://prntscr.com/8sumuz

OpenStudy (unicornpoopcookies):

http://openstudy.com/study#/updates/51d10538e4b08d0a48e3f951

OpenStudy (unicornpoopcookies):

check that maybe idk this math makes my head hurt lol

OpenStudy (kittiwitti1):

It usually happens as a glitch. If you reload I think it works. LaTex issues aha

OpenStudy (kittiwitti1):

@UnicornPoopCookies it's not the same equation asked for sorry

OpenStudy (unicornpoopcookies):

Oh well x)

OpenStudy (kittiwitti1):

I got csc wrong for #2 D:

Directrix (directrix):

The first question is asking where sine and cosine have the same value on the unit circle. Recall that the sine and cosine of an isosceles right triangle are the same because the adjacent and opposite legs are the same length. Sine t = Cosine t